OpenStudy (anonymous):

I will be forever gratefull and be your fan, if you could anwer my question!

OpenStudy (hoblos):

you dont have to... we are glad to help but where is the question??

OpenStudy (anonymous):

how do I solve: Find the volume of the cone z=sqrt(x^2+y^2) and inside the sphere x^2+y^2+z^2=a^2???

hero (hero):

How come your volume of cone doesn't include pi ?

OpenStudy (anonymous):

In polar coordinates, z=r.

OpenStudy (anonymous):

If it fits inside the sphere, r^2<a^2/2

OpenStudy (anonymous):

Since r=z, the volume of an differential disk in the cone is pi z^2 dz.

OpenStudy (anonymous):

Integrate from zero to a/sqrt (2) to get the answer.

OpenStudy (anonymous):

okay. i will try. thanks alot :)

OpenStudy (anonymous):

It might be close to\[\pi \int\limits_{0}^{a/\sqrt{2}}z^2dz=\pi(a/\sqrt{2}) ^{3}/3\]

OpenStudy (anonymous):

or not.... Hope I helped.

OpenStudy (anonymous):

you did. thanks mate :)